OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $3.95. If the number of dimes was doubled and the number of nickels was increased by 21, the value of the coins would be $8.40. How many nickels and dimes does he have?

OpenStudy (anonymous):

Let n be the number of nickles and d be the number of dimes. Then one equation is\[.05n+.10d=3.95\]and a second equation is\[.05(n+21)+.10*2d=8.40\]Solving this system will give the answer

OpenStudy (anonymous):

can someone please help Thanks!

OpenStudy (anonymous):

He has enough to buy a happy meal.

OpenStudy (curry):

with wat do iu need help

OpenStudy (anonymous):

help with solving

OpenStudy (curry):

post

OpenStudy (anonymous):

Multiply the first equation by 2 to get\[.1n+.2d=7.9\]and then solve it for .2d\[.2d=7.9-.1n\](continued)

OpenStudy (anonymous):

Distribute the .05 in the second equation\[.05n+1.05+.2d=8.4\]now substitute 7.9-.1n (from the last line in my previous post) in for the .2d in the equation above\[.05n+1.05+7.9-.1n=8.4\]This gives you an equation with a single variable, n. Solve for n. (continued)

OpenStudy (anonymous):

N = Nickels, D = Dimes So....N + D = 3.95 If he has 21 Nickels that means he has $1.05 So N = 1.05 1.05 + D = 8.40 D = 8.40 - 1.05 D = $7.35 Divide 7.35 by 10, hes got about 73 dimes

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Thanks gylvx Yes

OpenStudy (anonymous):

mandolino hes a kid, don't give him university equations lol

OpenStudy (anonymous):

it scares the pellet out of people that aren't math majors :D

OpenStudy (anonymous):

wait i didnt read the question properly

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

\[.05n-.1n=8.4-7.9-1.05\]\[-.05n=-.55\]\[n=\frac{.55}{.05}=11\]This is the number of nickles. To get the number of dimes substittute this value for n in any of the above equations\[.2d=7.9-.1n\]or\[.2d=7.9-.1(11)\]\[d=\frac{7.9-.1(11)}{.2}=34\]This the number of dimes.

OpenStudy (anonymous):

joy what grade are you in?

OpenStudy (anonymous):

Thank you Mandolino...:)

OpenStudy (anonymous):

I wish I would have kept my other medals..

OpenStudy (anonymous):

how do you give medals?

OpenStudy (anonymous):

@joy, You know, another way that might not seem so complicated would be to multiply each equation through by 100 to convert all the decimals to whole numbers. The process from there could be the same, but the arithmetic might seem easier. You are welcome for the help.

OpenStudy (anonymous):

Thank you so much again...great explanation..I got it now